# Multi-scale Graph Autoregressive Modeling: Molecular Property Prediction via Next Token Prediction

### v2

@@ 1 Introduction
Molecular property prediction is a core challenge in drug discovery (Wu et al., 2018) and has increasingly become a primary benchmark for “Foundation Models (FM) for Science.” (Khan et al., 2025) Within FM paradigm, the training recipe has standardized: decoder-only Transformers trained via next-token prediction (NTP) serve as the dominant engine for scale and generalization (Wang et al., 2024; Xia et al., 2025; Zhang et al., 2025). Consequently, as the core components of these training stacks become heavily optimized, the pragmatic research focus shifts away from bespoke backbone re-design toward optimizing the input representation to fit this proven architecture (Touvron et al., 2023). However, a critical gap remains: current NTP-based molecular models, which largely rely on 1D SMILES strings, fail to explicitly capture graph topology and consequently trail behind specialist graph models in predictive accuracy (Xia et al., 2025).
@@
To bridge this gap, we propose Connection-Aware Motif Sequencing (CamS), a tokenizer-level interface that makes molecular graphs directly learnable by standard decoder-only NTP through a three-stage design. First, to ensure information efficiency while preserving semantics, we adapt byte-pair encoding (BPE)-style mining on molecular graphs (Shibata et al., 1999; Geng et al., 2023; Shen and Poczos, ). This produces motif vocabulary with frequent merge operation and introduces a tunable molecular scale that allows direct granularity control. To ensure faithful encoding of fine-grained chemical detail, we further integrate a Single-Atom Vocabulary Closure (SAVC) mechanism. Second, to make the graph autoregressive-ready, we partition the molecule into a BPEGraph at a specific scale and serialize it via scaffold-rooted breadth-first search (BFS)  (Bundy and Wallen, 1984). This produces a CamS subsequence with a stable Intra-scale Order, moving from the global core to peripheral functional groups. Third, and most critically, to construct a hierarchical context, we concatenate subsequences from fine to coarse. This strategy circumvents the trade-off between scales and establishes an Inter-scale Order in the final CamS sequence, enabling high-level motifs to be predicted by conditioning on dense, uncorrupted fine-scale evidence. Collectively, these steps unify substructure compression and serialization through a dual-causal ordering—sequencing motifs within each scale and stacking scales hierarchically—to enable the model to comprehend global scaffolds based on explicit, uncorrupted local structural evidence.
@@
Built on CamS, we instantiate CamS-LLaMA by pre-training a native decoder-only backbone with standard NTP on CamS sequences, keeping the architecture and objective unchanged (Touvron et al., 2023). To keep attribution controlled, we inject only a lightweight molecular fingerprint (Cereto-Massagué et al., 2015) prior during downstream fine-tuning, while keeping pre-training purely sequence/structure-driven. Across MoleculeNet (Wu et al., 2018) and the activity-cliff stress test MoleculeACE (Van Tilborg et al., 2022), CamS-LLaMA achieves state-of-the-art performance at a comparable model scale to strong graph self-supervised learning baselines, despite relying on substantially weaker priors (Li et al., 2023). Beyond accuracy, we provide mechanism-level evidence. Targeted ablation studies confirm that multi-scale concatenation is the key driver of performance, while interpretability analysis reveals that the causal serialization effectively drives attention toward cliff-determining structural differences.
@@
Contributions. (1) Methodological Framework: We introduce CamS, a unified graph-to-causal-sequence interface that resolves the conflict between graph topology preservation and scalable autoregressive training, enabling NTP to function as a structure-native objective. (2) Empirical Validation: We demonstrate that CamS-LLaMA achieves state-of-the-art performance on MoleculeNet and MoleculeACE benchmarks. Mechanism analysis reveals that the multi-scale causal context explicitly improves attention focus on subtle, activity-determining structural edits. (3) Implementation Recipe: We provide a reproducible pipeline—from the CamS-tokenizer to the CamS-LLaMA—establishing a robust baseline for applying standard foundation model architectures to molecular science.
@@ 2 Related Work
2.1 Molecular Property Prediction within Foundation Models Paradigm
Molecular property prediction is increasingly adopting a foundation model approach: large-scale Transformer pre-training followed by fine-tuning for diverse property endpoints (Awais et al., 2025; Xia et al., 2025). One approach adopts standard large language model (LLM) methodology by representing molecules as SMILES strings and pre-training decoder-only models via NTP (Xia et al., 2025; Cai et al., 2025). This leverages mature LLM infrastructure and enables straightforward large-scale training. However, SMILES is not structure-native, weakening topology-based signals. String-based edits often misalign with actual chemical changes, and performance consistently trails strong graph-native methods on property prediction tasks (Xia et al., 2025). Alternatively, graph-native foundation models directly process molecular graphs to preserve connectivity and local topology (You et al., 2021; Liu et al., ; You et al., 2020; Xu et al., 2021). High-performing methods typically employ encoder-only Graph Transformers (Shehzad et al., 2024) designed for property prediction. While effective, these models deviate from the vanilla decoder-only NTP approach that exhibits superior generalization and scalability (Li et al., 2023; Rong et al., 2020). More recent structural foundation models integrate graphs with autoregressive components but frequently depend on additional objectives (often incorporating 3D structure) and learned molecular encoders rather than pure tokenizer-level NTP (Kong et al., 2025; Lu et al., 2025; Zhang et al., 2025).
@@ 2.2 Molecular Tokenization Strategies
Within a foundation model paradigm, tokenizer design becomes a key interface for scaling LLMs. In NLP, BPE yields strong compression via simple frequency-based merges without training an extra tokenizer model. Molecular tokenization spans string-based and graph-based methods (Wang et al., 2019; Liu et al., 2023). String-based schemes operate on SMILES (Weininger, 1988) or related linearizations (Krenn et al., 2022), while graph-based schemes have evolved from basic node/edge-level units to fragment-level motifs (Liu et al., 2023). Some rule-based motif tokenizations rely on limited handcrafted chemistry rules (Zhang et al., 2021); triplet tokenizations, as used in KPGT (Knowledge-guided Pre-training of Graph Transformer), show that enriching a node token with only one extra atom can substantially boost property prediction (Li et al., 2023); ring-or-path-based (Wollschläger et al., 2024) and tree-based (Jin et al., 2018) schemes also expose substructures at different granularities; some other works learn tokenization via trainable rules (Sun et al., 2024) or by training a GNN as a tokenizer (Liu et al., 2023; Luo et al., ). Particularly noteworthy are data-driven motif mining approaches that adapt BPE idea to graphs to obtain reusable, multi-scale fragments with strong compression and minimal reliance on manually designed rules or additional trained models (Geng et al., 2023; Shen and Poczos, ; Kong et al., 2022). In addition, cross-level tokenization schemes, which combine motifs at different structural levels, also provide complementary design ideas (Ji et al., 2022; Chen et al., ).
@@ 3 Method
Our molecular representation must satisfy two requirements: (1) chemical fidelity, i.e., sensitivity to pharmacophore-level structures and their topology; and (2) autoregressive foundation model compatibility, i.e., a compressed interface naturally aligned with an NTP paradigm (Touvron et al., 2023). Motivated by this, we propose Connection-Aware Motif Sequencing (CamS), which segments molecular graphs into reusable frequent motifs and, crucially, serializes the motif graph into multi-scale connection-aware causal sequences compatible with NTP pre-training (Section 3.1). Built on CamS, CamS-LLaMA performs large-scale autoregressive pre-training and injects a only lightweight fingerprint prior during downstream fine-tuning (Section 3.2). Finally, we analyze the characteristics and advantages of multi-scale CamS autoregressive modeling by comparing it with Graphformer-style modeling (Section 3.3).
@@ 3.1 CamS-Tokenizer
The CamS-tokenizer framework operates in three logical phases to bridge graph topology and scalable NTP. (1) CamS Vocabulary Construction: we maintain a robust vocabulary via BPE-style mining and Single-Atom Vocabulary Closure (SAVC) (Section 3.1.1). (2) Per-Scale Encoding: molecules are partitioned into motif graphs (BPEGraph) and serialized into single-scale CamS subsequences via scaffold-anchored BFS (Section 3.1.2). (3) Cross-Scale Concatenation: subsequences are organized in a fine-to-coarse order to form the final CamS sequence, enabling multiscale autoregressive modeling (Section 3.1.3). Full details are in Appendix A.
3.1.1 Substructure Level: CamS Vocabulary
BPE-style Mining. Following (Geng et al., 2023; Shen and Poczos, ), we first learn an ordered list of merge operations $\mathcal{O}$ from co-occurrence statistics. We then apply $\mathcal{O}$ to the corpus to mine the resulting subgraphs. This yields the BPEMotif set (frequent subgraphs with connectivity patterns) and the base BPE-SAV (Single-Atom Vocabulary with connectivity patterns) naturally captured during mining.
Single-Atom Vocabulary Closure (SAVC). While generation operates within a closed vocabulary, predictive encoding encounters unseen atom states. Standard GraphBPE tokenizers greedily match largest motifs and recursively back off; if a specific atom-connectivity form is absent, it collapses to [UNK]. This erases chemical details (e.g., element type) and creates Causal Information Breakpoints in the autoregressive stream. To prevent this, SAVC constructs the BPE-SAV Back-off via two strategies: (1) enumerating connection-aware tokens for all typical valences, and (2) mapping rare atypical forms of element X to <X_AltForm> rather than <UNK>.
The final CamS Vocabulary $\Sigma$ is the union of BPEMotif, the complete SAV (BPE-SAV $\cup$ Back-off), and special tokens mapped to unique integers. This guarantees that every atom is covered by either a high-level motif or a precise single-atom descriptor (details in Appendix A.1).
3.1.2 Per-Scale Encoding
Given a vocabulary scale $s$ (determined by the iteration count of merge operations), encoding a molecule $M$ involves three stages to generate a single-scale CamS subsequence.
BPEGraph Construction. We apply the prefix of merge operations to $M$, inducing a partition of atoms into non-overlapping fragments. This forms the BPEGraph $G_{s}=(V_{s},E_{s})$, where nodes $V_{s}\subset\Sigma$ are connection-aware motifs/atoms and $E_{s}$ represents chemical bonds between them.
Serialization (Intra-scale Order). To adapt the graph for decoder-only NTP, we serialize $V_{s}$ into an ordered list via Breadth-First Search (BFS). We select the motif with the largest atom count (typically the core scaffold) as the root. This strategy prioritizes global backbone structure before local substitutions (Zhang et al., 2021), establishing a stable Center-to-Periphery causal order.
ID Extraction. Mapping the ordered nodes to their indices in $\Sigma$ yields the single-scale CamS subsequence $X^{(s)}=(x^{(s)}_{1},\dots,x^{(s)}_{L_{s}})$.
3.1.3 Cross-Scale Concatenation
@@
Single-scale tokenization faces a trade-off: coarse scales capture global scaffolds but offer low resolution, while fine scales preserve atomic details but fragment high-level structures (e.g., pharmacophores). CamS resolves this by constructing a multi-scale causal context. We tokenize the molecule at multiple scales (e.g., $s\in\{1\mathrm{K},7\mathrm{K},27\mathrm{K},67\mathrm{K}\}$) by slicing $\mathcal{O}$. The resulting subsequences are concatenated in a fixed Fine-to-Coarse order (Inter-scale Order) to form the final CamS sequence $\mathbf{X}$:
$\mathbf{X}=[$ $\,\texttt{[BOS]},X^{(1\mathrm{K})},\texttt{[CONCAT]},X^{(7\mathrm{K})},\texttt{[CONCAT]},$ (1)
$\,X^{(27\mathrm{K})},\texttt{[CONCAT]},X^{(67\mathrm{K})},\texttt{[EOS]}].$
@@
During training, this CamS token stream allows the model to leverage high-resolution local topology (fine scales) as a prefix to condition the prediction of global scaffolds (coarse scales). This effectively embeds bottom-up structural composition into the autoregressive objective without architectural modifications (analysis in Section 3.3).
@@ 3.2 CamS-LLaMA
Figure 1: The overall framework of CamS-LLaMA. (a) CamS-Tokenizer: Molecules are transformed into multi-scale causal sequences. Through Per-scale Encoding, motif graphs are serialized via Scaffold-rooted BFS (Intra-Scale Order) to establish CamS Subsequence, followed by Cross-Scale Concatenation to arrange subsequences into a final CamS Sequence from fine to coarse granularities (Inter-Scale Order). (b) Pre-training Pipeline: The model is pre-trained on CamS Token Lists using a standard LLaMA via NTP, learning to aggregate structural information. (c) Fine-tuning Pipeline: During downstream tasks, the pre-trained weights are transferred, and a lightweight fingerprint prior is injected (via projection, prepending in embedding layer and fusion in classifier or regressor) to take part in the prediction of specific properties.
CamS-LLaMA follows a representation-model co-design principle. Section 3.1 maps a molecular graph $G=(V,E)$ into a multi-scale, connection-aware causal sequence $X=(x_{1},\ldots,x_{T})$ (Eq. (1)) under vocabularies $\Sigma_{1\mathrm{K}}\subset\Sigma_{7\mathrm{K}}\subset\Sigma_{27\mathrm{K}}\subset\Sigma_{67\mathrm{K}}\triangleq\Sigma$, via BFS-ordered motif traversals $\mathcal{T}_{s}(G)$ and token mapping $x^{(s)}_{\ell}=\iota_{s}(\tau^{(s)}_{\ell})\in\Sigma_{s\mathrm{K}}\subseteq\Sigma$. The key point is that this graph-to-causal-sequence interface is intentionally designed to be transparent to a vanilla decoder-only Transformer: it exposes topology through a causal order and a fine-to-coarse hierarchy, so we can train with standard NTP without modifying the LLaMA backbone. Based on this interface, CamS-LLaMA follows the foundation model protocol: (1) self-supervised autoregressive pre-training on CamS sequences (figure 1 bottom right), and (2) supervised fine-tuning with a minimal structural prior (fingerprints) injected only at fine-tuning time (figure 1 bottom left).
3.2.1 Autoregressive Pre-training via NTP

### v3

@@ 1 Introduction
Molecular property prediction is a core challenge in drug discovery (Wu et al., 2018) and has increasingly become a primary benchmark for “Foundation Models (FMs) of science.” (Khan et al., 2025) Within FM paradigm, the training recipe has standardized: decoder-only Transformers trained via next-token prediction (NTP) serve as the dominant engine for scale and generalization (Wang et al., 2024; Xia et al., 2025; Zhang et al., 2025). Consequently, as the core components of these training stacks become heavily optimized, the pragmatic research focus shifts away from bespoke backbone re-design toward optimizing the input representation to fit this proven architecture (Touvron et al., 2023). However, a critical gap remains: current NTP-based molecular models, which largely rely on 1D SMILES strings, fail to explicitly capture graph topology and consequently trail behind specialist graph models in predictive accuracy (Xia et al., 2025).
@@
To bridge this gap, we propose Connection-Aware Motif Sequencing (CamS), a tokenizer-level interface that makes molecular graphs directly learnable by standard decoder-only NTP through a three-stage design. First, to ensure information efficiency while preserving semantics, we adapt byte-pair encoding (BPE)-style mining on molecular graphs (Shibata et al., 1999; Geng et al., 2023; Shen and Poczos, 2024). This produces motif vocabulary with frequent merge operation and introduces a tunable molecular scale that allows direct granularity control. To ensure faithful encoding of fine-grained chemical detail, we further integrate a Single-Atom Vocabulary Closure (SAVC) mechanism. Second, to make the graph autoregressive-ready, we partition the molecule into a BPEGraph at a specific scale and serialize it via scaffold-rooted breadth-first search (BFS)  (Bundy and Wallen, 1984). This produces a CamS subsequence with a stable Intra-scale Order, moving from the global core to peripheral functional groups. Third, and most critically, to construct a hierarchical context, we concatenate subsequences from fine to coarse. This strategy circumvents the trade-off between scales and establishes an Inter-scale Order in the final CamS sequence, enabling high-level motifs to be predicted by conditioning on dense, uncorrupted fine-scale evidence. Collectively, these steps unify substructure compression and serialization through a dual-causal ordering—sequencing motifs within each scale and stacking scales hierarchically—to enable the model to comprehend global scaffolds based on explicit, uncorrupted local structural evidence.
@@
Built on CamS, we instantiate CamS-LLaMA by pre-training a native decoder-only backbone with standard NTP on CamS sequences, keeping the architecture and objective unchanged (Touvron et al., 2023). To keep attribution controlled, we inject only a lightweight molecular fingerprint (Cereto-Massagué et al., 2015) prior during downstream fine-tuning, while keeping pre-training purely sequence/structure-driven. Across MoleculeNet (Wu et al., 2018) and the activity-cliff stress test MoleculeACE (Van Tilborg et al., 2022), CamS-LLaMA achieves state-of-the-art (SOTA) performance at a comparable model scale to strong graph self-supervised learning baselines, despite relying on substantially weaker priors (Li et al., 2023). Beyond accuracy, we provide mechanism-level evidence. Targeted ablation studies confirm that multi-scale concatenation is the key driver of performance, while interpretability analysis reveals that the causal serialization effectively drives attention toward cliff-determining structural differences.
@@
Contributions. (1) Methodological Framework: We introduce CamS, a unified graph-to-causal-sequence interface that resolves the conflict between graph topology preservation and scalable autoregressive (AR) training, enabling NTP to function as a structure-native objective. (2) Empirical Validation: We demonstrate that CamS-LLaMA achieves SOTA performance on MoleculeNet and MoleculeACE benchmarks. Mechanism analysis reveals that the multi-scale causal context explicitly improves attention focus on subtle, activity-determining structural edits. (3) Implementation Recipe: We provide a reproducible pipeline—from the CamS-tokenizer to the CamS-LLaMA—establishing a robust baseline for applying standard AR FM architectures to molecular science.
@@ 2 Related Work
2.1 Molecular Property Prediction via FMs
Molecular property prediction is increasingly adopting a FM approach: large-scale Transformer pre-training followed by fine-tuning for diverse property endpoints (Awais et al., 2025; Xia et al., 2025). One approach adopts standard large language model (LLM) methodology by representing molecules as SMILES strings and pre-training decoder-only models via NTP (Xia et al., 2025; Cai et al., 2025). This leverages mature LLM infrastructure and enables straightforward large-scale training. However, SMILES is not structure-native, weakening topology-based signals. String-based edits often misalign with actual chemical changes, and performance consistently trails strong graph-native methods on property prediction tasks (Xia et al., 2025). Alternatively, graph-native FMs directly process molecular graphs to preserve connectivity and local topology (You et al., 2021; Liu et al., 2022; You et al., 2020; Xu et al., 2021). High-performing methods typically employ encoder-only Graph Transformers (Shehzad et al., 2024) designed for property prediction. While effective, these models deviate from the vanilla decoder-only NTP approach that exhibits superior generalization and scalability (Li et al., 2023; Rong et al., 2020). More recent structural FMs integrate graphs with AR components but frequently depend on additional objectives (often incorporating 3D structure) and learned molecular encoders rather than pure tokenizer-level NTP (Kong et al., 2025; Lu et al., 2025; Zhang et al., 2025).
@@ 2.2 Molecular Tokenization Strategies
Within a FM paradigm, tokenizer design becomes a key interface for scaling backbones. String-based schemes operate on SMILES (Weininger, 1988; Wang et al., 2019) or SELFIES (Krenn et al., 2022). Graph-based approaches employ diverse substructure definitions: rule-based tokenizations rely on limited handcrafted chemistry rules (Zhang et al., 2021); triplet tokenizations, as used in KPGT (Knowledge-guided Pre-training of Graph Transformer), show that enriching a node token with only one extra atom can substantially boost property prediction (Li et al., 2023); ring-or-path-based (Wollschläger et al., 2024) and tree-based (Jin et al., 2018) schemes also expose substructures at different granularities; and other works learn tokenization via trainable rules (Sun et al., 2024) or graph neural networks (Liu et al., 2023; Luo et al., 2025). Particularly noteworthy are data-driven approaches that adapt BPE idea to graphs to obtain reusable fragments with strong compression and minimal reliance on rules or additional trained models (Geng et al., 2023; Shen and Poczos, 2024; Kong et al., 2022). In addition, cross-scale schemes, which combine motifs at different structural levels, also provide complementary design ideas (Ji et al., 2022; Chen et al., 2025).
@@ 3 Method
To align chemical fidelity with AR compatibility, we propose Connection-Aware Motif Sequencing (CamS). This interface serializes molecular graphs into multi-scale, causal sequences naturally consumable by decoder-only models (Section 3.1). Instantiated on this representation, CamS-LLaMA performs standard NTP pre-training, injecting a lightweight fingerprint prior only during downstream fine-tuning (Section 3.2). We conclude by theoretically contrasting this CamS-LLaMA framework with Graph Transformers paradigm (Section 3.3).
@@ 3.1 CamS-Tokenizer
The CamS-tokenizer framework bridges graph topology and scalable NTP through data-driven fragment compression at the substructure level and hierarchical serialization (both intra- and inter-scale) at the molecule level. Full implementation details are provided in Appendix A.
CamS Vocabulary Construction. On the substructure level, we adopt a data-driven strategy grounded in the philosophy that compression implies understanding. Unlike rigid manual rules, we apply BPE-style Mining to capture statistically significant chemical semantics. Following (Geng et al., 2023; Shen and Poczos, 2024), we first learn an ordered list of merge operations $\mathcal{O}$, sorted by descending frequency based on co-occurrence statistics. Since BPE merges are iterative, later operations in $\mathcal{O}$ (lower frequency) build upon earlier ones to combine smaller fragments into larger complex structures. Thus, the list order represents a trajectory from fine-grained local substructures to coarse-grained global scaffolds. Applying the full list $\mathcal{O}$ yields the BPEMotif set (frequent subgraphs vocabulary) and the base BPE-SAV (single atoms vocabulary) naturally captured during mining.
Crucially, distinct from prior works, we introduce Single-Atom Vocabulary Closure (SAVC) to address a critical flaw: while generation operates within a closed vocabulary, predictive encoding encounters unseen atom states. Standard tokenizers greedily back off to [UNK] when specific atom-connectivity forms are absent, erasing element details and creating “Causal Information Breakpoints” in the AR stream. To prevent this, SAVC constructs the BPE-SAV Back-off by: (1) enumerating connection-aware tokens for all typical valences, and (2) mapping rare atypical forms of element X to [X_AltForm] rather than [UNK]. The final CamS Vocabulary $\Sigma$ is the rigorous union of BPEMotif, the complete SAV (BPE-SAV $\cup$ Back-off), and special tokens. This design not only guarantees that every atom is covered by either a high-level motif or a precise single-atom descriptor but also provides a pivotal manipulable dimension for downstream representation: the Motif Scale $s$. Formally, we define the Motif Scale $s$ as the effective vocabulary size induced by a specific granularity level. Each scale $s$ corresponds to a specific prefix length $k$ of the merge list $\mathcal{O}$. A small $s$ (requiring fewer merges) activates only high-frequency motifs, keeping the graph fragmented (Fine Scale); conversely, a large $s$ (using more merges) includes rare, late-stage operations that aggregate these fragments into larger scaffolds (Coarse Scale).
Per-Scale Encoding. On the molecule level, utilizing the constructed CamS Vocabulary $\Sigma$, we can transform raw atom nodes into discrete substructure nodes containing rich connectivity, contextual compression and controllable scale. Specifically, given a molecular graph $G_{\text{mol}}$ and a target Motif Scale $s$ (associated with the operation prefix $\mathcal{O}_{\leq k}$), we apply the corresponding $k$ merge operations to $G_{\text{mol}}$. This process induces a partition of atoms into non-overlapping fragments, forming the BPEGraph $G_{s}=(V_{s},E_{s})$. Here, nodes $V_{s}\subset\Sigma$ denote connection-aware motifs/atoms (ranging from fine fragments to coarse scaffolds depending on $s$) and $E_{s}$ represents the chemical bonds preserving connectivity between them. Subsequently, a key challenge remains in how the NTP perceives this high-quality BPEGraph: the definition of causal order. While graphs permit various traversals (e.g., random walks (Lovász, 1993)), we establish a deterministic Intra-scale Order. We select the motif with the largest atom count (typically the core scaffold) as the root and serialize $V_{s}$ into an ordered list via Scaffold-Rooted BFS. This strategy prioritizes global backbone structure before local substitutions (Zhang et al., 2021), establishing a stable Center-to-Periphery causal order. Finally, we apply ID Extraction to map ordered nodes to their indices in $\Sigma$, yielding the single-scale CamS subsequence $X^{(s)}=(x^{(s)}_{1},\dots,x^{(s)}_{L_{s}})$.
@@
Cross-Scale Concatenation. The manipulable Motif Scale $s$ introduces an inherent resolution trade-off: coarse scales capture global scaffolds but offer low resolution, while fine scales preserve atomic details but fragment high-level pharmacophores. CamS resolves this by constructing a multi-scale causal context. We tokenize the molecule at a sequence of $M$ increasing scales $\mathcal{S}=\{s_{1},s_{2},\dots,s_{M}\}$. The resulting CamS Subsequences $\{X^{(s)}\mid s\in\mathcal{S}\}$ are concatenated in a fine-to-coarse order (i.e., the Inter-Scale Order) to form the final CamS sequence $\mathbf{X}$:
$\mathbf{X}=[$ $\,\texttt{[BOS]},X^{(s_{1})},\texttt{[CONCAT]},X^{(s_{2})},\dots,$ (1)
$\,\texttt{[CONCAT]},X^{(s_{M})},\texttt{[EOS]}].$
@@
During training, this CamS token stream allows the model to leverage high-resolution local topology (fine scales) as a prefix to condition the prediction of global scaffolds (coarse scales). This effectively embeds bottom-up structural composition into the AR objective without architectural modifications (analysis in Section 3.3).
Figure 1: The overall framework of CamS. (a) CamS-Tokenizer: Transforming molecules into causal sequences via the manipulable Motif Scale $s$. First, Per-scale Encoding applies merge operations to construct motif graphs, which are serialized via Scaffold-rooted BFS to establish Intra-Scale Order. Subsequently, Cross-Scale Concatenation arranges these views from fine ($s=1\mathrm{K}$) to coarse ($s=67\mathrm{K}$) to establish Inter-Scale Order. The resultant CamS sequence serves as a native token stream for a vanilla LLaMA backbone. (b) CamS-LLaMA: The model is pre-trained on the resultant CamS Token Stream via standard NTP. Pre-trained weights are transferred, and a lightweight fingerprint prior is injected during fine-tuning.
@@ 3.2 CamS-LLaMA
Following a representation-model co-design principle, CamS-LLaMA leverages the graph-to-sequence interface established in Section 3.1 to enable standard AR modeling on molecular graphs. As illustrated in Figure 1(a), the multi-scale CamS sequence acts as a transparent bridge, exposing topology to a vanilla LLaMA backbone without architectural modifications. This facilitates a scalable two-stage FM pipeline (Figure 1(b)): (1) self-supervised pre-training via NTP to capture structural generative laws, and (2) property-guided fine-tuning with a controlled fingerprint prior.